OpenStudy (anonymous):

show the algebraic inequality. show work. Tammy has the following scores on her business law test 92,99,84,91. what score must she earn in order to an A average. and A at her univ starts at 90

OpenStudy (gw2011):

(92+99+84+91+x)/5>=90 Where the symbol >= represents greater than or equal to and x represents the score needed 92+99+84+91+x>=(90)(5) 92+99+84+91+x>=450 Now add the numbers on the left side to get: 366+x>=450 x>=450-366 x>=84 Therefore, 84 is the lowest score needed in order to get an A average.